# ULPT: Prompt Tuning with Ultra-Low-Dimensional Optimization

### v1

@@ Abstract
Large language models achieve state-of-the-art performance but are costly to fine-tune due to their size. Parameter-efficient fine-tuning methods, such as prompt tuning, address this by reducing trainable parameters while maintaining strong performance. However, prior methods tie prompt embeddings to the model’s dimensionality, which may not scale well with larger LLMs and more customized LLMs. In this paper, we propose Ultra-Low-dimensional Prompt Tuning (ULPT), which optimizes prompts in a low-dimensional space (e.g., 2D) and use a random but frozen matrix for the up-projection. To enhance alignment, we introduce learnable shift and scale embeddings. ULPT drastically reduces the trainable parameters, e.g., 2D only using 2% parameters compared with vanilla prompt tuning while retaining most of the performance across $21$ NLP tasks. Our theoretical analysis shows that random projections can capture high-rank structures effectively, and experimental results demonstrate ULPT’s competitive performance over existing parameter-efficient methods. [^1] Our code is available at https://github.com/MANGA-UOFA/ULPT
@@ 1 Introduction
Prompt tuning introduces learnable prompt embeddings exclusively in the input layer of the model (Lester et al., 2021; Liu et al., 2024), automating prompt engineering by gradient descent to guide the frozen LLM in producing task-specific outputs (Petrov et al., 2024b, a). By contrast, LoRA modifies the model by injecting low-rank weight matrices into its layers, causing the number of trainable parameters to scale with model’s depth (Hu et al., 2022). Given that LLMs encode substantial knowledge during pretraining (Brown et al., 2020; Kojima et al., 2022) and that both in-context learning and expertly crafted prompts can achieve remarkable results (Wei et al., 2022b; Dong et al., 2024), prompt tuning offers a more efficient and effective alternative to LoRA in many scenarios (Shi and Lipani, 2024).
@@
Despite its advantages, most existing prompt tuning approaches couple the dimensionality of prompt embeddings with the hidden size of the model (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2022; Choi et al., 2023; Razdaibiedina et al., 2023). As the size of the model increases, the dimensionality of the prompt embedding space also increases Raffel et al. (2020); Touvron et al. (2023). This scaling leads to unnecessary complexity, as full dimensionality is often not required for task adaptation (Aghajanyan et al., 2021; Qin et al., 2022). Consequently, optimizing in this expanded space becomes inefficient in parameter’s usage and may also increase the risk of overfitting, especially for less complex tasks or with limited training data.
@@
In this paper, we propose Ultra-Low-Dimensional Prompt Tuning (ULPT), a method that decouples the prompt/model dimensions and enables prompt tuning in an ultra-low-dimensional space (e.g., 2D). A naïve attempt is to jointly optimize the ultra-low-dimensional embeddings with an up-projection matrix (Xiao et al., 2023; Guo et al., 2024), but the learnable up-projection matrix may result in more trainable parameters than vanilla prompt tuning, therefore offsetting the gains in parameter efficiency. As shown in Figure 1a, our ULPT eliminates this overhead by using a random but frozen matrix for the up-projection. We further introduce learnable shift and scale embedding vectors to better align the up-projected embeddings and the model’s prompt space (Wu et al., 2024c). In addition, we provide theoretical analysis for our ULPT, which not only proves its convergence but also shows that a low-dimensional space with random projection can effectively approximate high-rank information.
@@
Figure 1: Overview of our approach. (a) ULPT up-projects ultra-low-dimensional embeddings with a random but fixed matrix, followed by a learnable alignment mechanism shared across all up-projected embeddings. (b) ULPT can significantly reduce parameters usage for LLM customizations.
The ultra-low-dimensional nature of our ULPT is particularly suitable for scenarios requiring massive LLM customizations Mangrulkar et al. (2022) and continual learning Wang et al. (2022), as shown in Figure 1b. For example, in a typical prompt tuning approach, each task might require 100K real-valued parameters, which can add up significantly when scaling to millions of customized LLMs or tasks. By contrast, our ULPT with 2D prompt embeddings can reduce this to just 2K parameters per task, presenting a dramatic parameter saving to just 2% of the original usage.
We conducted experiments across $21$ NLP datasets to evaluate our ULPT. The results demonstrate that ULPT can extend a few-token vanilla prompt tuning setup to a $100$-token configuration without increasing the number of trainable parameters, while matching the performance of a fully parameterized $100$-token prompt tuning setup. With appropriate settings of the dimension, ULPT outperforms existing prompt tuning-based methods, while requiring much fewer trainable parameters.
@@
  • We introduce ULPT (Ultra-Low-Dimensional Prompt Tuning), which optimizes prompts in a low-dimensional space with a random up-projection and learnable shift and scale vectors, drastically reduces trainable parameters while maintaining performance.
  • We provide theoretical analysis showing ULPT’s ability to capture high-rank structures effectively and ensure convergence.
  • Across $21$ NLP tasks, ULPT delivers comparable performance to vanilla prompt tuning while reducing trainable parameters 2% of the original usage. When scaling to higher dimensions, it outperforms existing prompt tuning-based methods with much fewer trainable parameters.
@@ 2 Related Work
Parameter-efficient fine-tuning. With the rapid growth of pretrained neural networks, researchers have investigated parameter-efficient fine-tuning methods that update only a small set of parameters while maintaining high performance. One straight way is to tune specific components of the model. For example, BitFit Ben Zaken et al. (2022) updates only the bias terms, and LayerNorm tuning Zhao et al. (2024) only trains the layer-norm parameters. Another line of work involves introducing and training small, task-specific non-linear modules, such as Adapters Houlsby et al. (2019) and AdapterDrop Rücklé et al. (2021). Other methods steer the activation representations either globally or locally Wu et al. (2024b); Yin et al. (2024).
@@
Two other prominent paradigms are low-rank adaptation (LoRA; Hu et al., 2022) and prompt tuning methods Lester et al. (2021), which are more related to our work. They will be further elaborated in the subsequent sections.
@@
Low-rank adaptation. Hu et al. (2022) assume that weight updates can be approximated by low-rank matrices and propose a low-rank adaptation (LoRA) method for fine-tuning a model. Building upon this foundational work, many extensions have been developed to enhance LoRA’s performance. For example, ReLoRA Lialin et al. (2024) iteratively trains and merges low-rank adapters to achieve high-rank updates. Hayou et al. (2024) propose learning low-rank matrices with different learning rates. Wu et al. (2024a) explore training a mixture of LoRA modules and leverage dynamic routing mechanisms for different task distributions or domains.
However, for large models, LoRA still requires a considerable number of trainable parameters. To address this limitation, several works have explored the use of random projection Bingham and Mannila (2001) to further improve parameter efficiency. For example, Flora Hao et al. (2024) updates the pretrained matrices with randomly projected gradients, while VeRA Kopiczko et al. (2024) uses random projections combined with two trainable scaling vectors to approximate each update matrix.
@@
Prompt tuning. Shin et al. (2020) introduce the concept of learning prompt tokens to elicit knowledge from LLMs. Subsequently, Lester et al. (2021) extend this idea to continuous prompt tuning, where prompt embeddings are optimized through gradient descent while keeping the LLM frozen. Building on this, Li and Liang (2021) further generalize prompt embeddings to a multi-layer setting. Razdaibiedina et al. (2023) re-parameterize prompt tuning by incorporating a feed-forward neural network with residual connections. Shi and Lipani (2024) observe that redistributing parameters to learn offsets for input token embeddings can enhance performance. On the other hand, multi-task prompt tuning has been explored, where the learned prompt parameters are reused across different tasks Wang et al. (2023). Closely with our work, Xiao et al. (2023) decompose the prompt embedding matrix into two low-rank components: a low-dimensional prompt matrix and a learnable up-projection matrix. By contrast, our ULPT method freezes the up-projection matrix, so that we are able to achieve high performance with much fewer trainable parameters, supported by random-projection theory Bingham and Mannila (2001). Overall, our approach is able to function well with an ultra-low dimension, making it practical to customize millions of LLMs and perform continual learning in an ever-changing environment.
@@ 3.1 Problem Formulation
Figure 2: Distribution of prompt embedding values over 100 prompt tokens. We randomly selected $20$ dimensions from the original prompt embeddings, which have 768 dimensions as in the T5-base model. The mean, 25/75 percentiles, and min/max are shown for the embedding values learned in the CoLA and SST-2 tasks (details explained in §4.1).
3.2 Our Ultra-Low-Dimensional Prompt Tuning
@@
A naïve implementation of this decomposition treats both ${\bm{Z}}$ and ${\bm{P}}$ as learnable parameters Xiao et al. (2023), which reduces the number of trainable parameters to $nr+rd$. This, unfortunately, scales poorly for larger models, as an $r\times d$ up-projection matrix should be learned and stored, which undermines the savings of learnable parameters.
@@
In this way, we completely eliminate the need for storing the up-project matrix. This not only largely reduces the learnable parameters from $nr+rd$ to $nr$ (plus one extra random seed), but also combats overfitting especially when the fine-tuning dataset is small.
In our pilot study, we observe that typical prompt embeddings ${\bm{E}}$, even without low-rank treatment, exhibit significant variation across different dimensions, as shown in Figure 2. These variations may cause difficulty during training, therefore, we further introduce a learnable shift embedding $\bm{s}\in\mathbb{R}^{d}$ and a learnable scale embedding $\bm{b}\in\mathbb{R}^{d}$ to adjust the projected embeddings to ensure better alignment with the varying distributions across dimensions. Notice that the shift and scale embeddings are shared across different prompt token positions, but may vary for different tasks.
@@
where $z_{ik}$ and $\tilde{p}_{kj}$ are an entry in ${\bm{Z}}$ and $\tilde{\bm{P}}$ matrices, respectively; $s_{j}$ and $b_{j}$ are an entry in $\bm{s}$ and $\bm{j}$ vectors, respectively.
@@ 3.3 Theoretical Analyses
theoremexpr Let ${\bm{e}}_{1},\cdots,{\bm{e}}_{n}\in\mathbb{R}^{d}$ be the embedding vectors in the high-dimensional space. Let ${\bm{P}}\in{\mathbb{R}}^{r\times d}$ be a random projection matrix with each element $p_{i,j}\sim\mathcal{N}(0,1/r)$. There exists a set of low-dimensional vectors ${\bm{z}}_{1},\cdots,{\bm{z}}_{n}\in\mathbb{R}^{r}$ such that with confidence at least $1-\delta$ we have
@@
for all $i,j\in[n]$, as long as $r\geq 2c\epsilon^{-2}\log(2n/\delta)$.
@@
In essence, our theorem asserts that there exists a set of low-dimensional vectors such that the pair-wise $L^{2}$ distances of the original high-dimensional vectors are preserved for all $(i,j)$ pairs. More importantly, the projected dimension $r$ only grows logarithmically with the original dimension $n$, demonstrating a favorable property of scaling. It should be noted that, although our theorem uses $L^{2}$ as the metric, it can be easily extended to the dot-product metric as well by noticing that $\|{\bm{x}}-{\bm{y}}\|^{2}=\|{\bm{x}}\|^{2}+\|{\bm{y}}\|^{2}-2{\bm{x}}\cdot{\bm{y}}$.
@@
The above theorem shows the existence of an expressive low-dimensional space. We assert in the following theorem that, given a random up-projection matrix, the optimal low-dimensional embeddings can be learned by gradient descent under mild assumptions.
@@
Theorem 3.3 shows that, even with the naïve gradient descent, the fixed random matrix ${\bm{P}}$ does not hinder the optimization procedure. By combining Theorem 1, we theoretically justifies our overall practice of ULPT.
4 Experiments
4.1 Experimental Settings
Datasets. We evaluate the proposed ULPT method across $21$ NLP tasks following prior work Asai et al. (2022); Wang et al. (2023); Shi and Lipani (2024). Those tasks are grouped into 4 categories: (1) GLUE Wang et al. (2018) is a benchmark suite consisting of various language understanding tasks, such as MNLI Williams et al. (2018), QQP Wang et al. (2018), QNLI Demszky et al. (2018), SST-2 Socher et al. (2013), STS-B Cer et al. (2017), MRPC Dolan and Brockett (2005), RTE Giampiccolo et al. (2007) and CoLA Warstadt et al. (2019). (2) SuperGLUE Wang et al. (2019) extends GLUE with more challenging tasks with limited training data, consisting of MultiRC Khashabi et al. (2018), BoolQ Clark et al. (2019), WiC Pilehvar and Camacho-Collados (2019), WSC Levesque et al. (2012), and CB De Marneffe et al. (2019). (3) The MRQA 2019 Shared Tasks Fisch et al. (2019) are a set of QA tasks to test LLM generation capabilities, consisting of Natural Questions Kwiatkowski et al. (2019), HotpotQA Yang et al. (2018), SearchQA Dunn et al. (2017) and NewsQA Trischler et al. (2017). (4) Other tasks beyond the above test suites are also considered, including WinoGrande Sakaguchi et al. (2021), Yelp-2 Zhang et al. (2015), SciTail Khot et al. (2018), and PAWS-Wiki Zhang et al. (2019). Further details on these datasets are provided in Table 5 in Appendix B.
Baselines. We compare our ULPT against a wide range of baselines to demonstrate its effectiveness and parameter efficiency. First, we evaluate against full-model fine-tuning, which optimizes all model parameters for downstream task adaptation, serving as a strong but parameter-intensive baseline. Second, we include state-of-the-art parameter-efficient methods such as Adapter Houlsby et al. (2019), AdapterDrop Rücklé et al. (2021), BitFit Ben Zaken et al. (2022), HyperFormer Karimi Mahabadi et al. (2021), HyperDecoder Ivison and Peters (2022), LoRA Hu et al. (2022), and Ladder Side-Tuning (LST; Sung et al., 2022). Third, we compare ULPT with vanilla prompt tuning (PT) and its variants, including DePT Shi and Lipani (2024), which learns offsets to the input token embeddings while using a separate learning rate for the prompt embeddings, and DPT Xiao et al. (2023), which is closely related to ULPT and decomposes prompt embeddings into low-rank matrices. Finally, we compare ULPT with transfer or multi-task learning methods, including SPoT Vu et al. (2022), ATTEMPT Asai et al. (2022), and MPT Wang et al. (2023).
Implementation details. In our pilot study (Figure 2), we perform vanilla prompt tuning on the T5-base model Raffel et al. (2020) with CoLA and SST-2, using $n=100$ prompt embeddings, each having a dimensionality of $d=768$. We randomly select $20$ dimensions and report the mean, $25$th/$75$th percentiles, and the minimum/maximum values for each dimension.
In our main experiment, we use the T5-base model with $d=768$. Consistent with prior work Shi and Lipani (2024); Xiao et al. (2023), we set the number of prompt tokens $n=100$ for the prompt embeddings ${\bm{Q}}\in{\mathbb{R}}^{n\times r}$ of our ULPT. For the rank $r$, we evaluate four configurations: $r=2$, $16$, $64$, and $256$, ranging from an ultra-low-dimensional setup to a more expressive configuration of $1/3$ of the original prompt dimension. All experiments use a batch size of $16$ and a default learning rate of $6e{-1}$ with AdamW. The learning rate follows a linear schedule, warming up for $500$ steps and then decaying linearly to 00. We set a maximum sequence length of $256$ for most tasks, except for SuperGLUE-MultiRC being $348$ and MRQA being $512$. ULPT is trained on all tasks for up to $100,000$ steps. Performance is evaluated every $1,000$ steps, with the best checkpoint selected based on the validation set.
In our analysis, T5-small ($d=512$) and T5-large model ($d=1024$) are considered to evaluate the generality of ULPT across different model sizes. We also vary the number of prompt tokens from $10$ to $100$ under different rank configurations. Further details are provided in §4.3.
Table 1: Performance on GLUE and SuperGLUE benchmarks based on the T5-base model. We report standard evaluation metrics, namely, Pearson correlation for STS-B, F1 for MultiRC, and accuracy for other tasks. We replicate prompt tuning (PT; Lester et al., 2021) and DPT Xiao et al. (2023) using their default configurations. Our replicated PT results slightly exceed those reported in previous studies. All other baseline results are directly sourced from Shi and Lipani (2024). The suggested rank for DPT is r=10 based on Xiao et al. (2023); we replicate DPT with r=64 for controlled comparison with our ULPT. tTransfer learning methods. mMulti-task learning methods, whose “#param/task” scores are calculated based on the GLUE benchmark.

### v2

# Ultra-Low-Dimensional Prompt Tuning via Random Projection

@@ Abstract
Large language models achieve state-of-the-art performance but are increasingly costly to fine-tune. Prompt tuning is a parameter-efficient fine-tuning method that addresses parameter-efficiency by learning prompt embeddings, but these embeddings are typically tied to the model’s hidden dimensionality, limiting parameter saving. In this paper, we propose Ultra-Low-dimensional Prompt Tuning (ULPT), a simple yet effective method that optimizes prompts in a low-dimensional space (e.g., 2D) and uses a frozen random matrix for up-projection. ULPT can achieve 98% reduction in the training parameters compared to vanilla prompt tuning while preserving performance. Our extensive experiments across over 20 NLP tasks demonstrate that ULPT consistently outperforms recent parameter-efficient tuning methods using significantly fewer parameters, making it well-suited as a storage-efficient framework for massive LLM customization. [^1] Our code is available at https://github.com/MANGA-UOFA/ULPT
Ultra-Low-Dimensional Prompt Tuning via Random Projection
Zijun Wu, Yongchang Hao, Lili Mou Dept. Computing Science & Alberta Machine Intelligence Institute (Amii), University of Alberta Canada CIFAR AI Chair {zijun4,yongcha1}@ualberta.ca, doublepower.mou@gmail.com
@@ 1 Introduction
Prompt tuning introduces learnable prompt embeddings exclusively in the input layer of the model (Lester et al., 2021; Liu et al., 2024a), automating prompt engineering by gradient descent to guide the frozen LLM in producing task-specific outputs (Petrov et al., 2024b, a). By contrast, LoRA modifies the model by injecting low-rank weight matrices into its layers, causing the number of trainable parameters to scale with model’s depth (Hu et al., 2022). Given that LLMs have encoded substantial knowledge during pretraining (Brown et al., 2020; Kojima et al., 2022) and that both in-context learning and expertly crafted prompts can achieve remarkable results (Wei et al., 2022b; Dong et al., 2024), prompt tuning offers a more efficient and effective alternative to LoRA in many scenarios (Shi and Lipani, 2024).
@@
Despite these advantages, a key limitation remains: prompt embeddings are typically constrained to match the model’s hidden dimensionality (Lester et al., 2021; Li and Liang, 2021; Liu et al., 2022b; Choi et al., 2023; Razdaibiedina et al., 2023). As the size of the model increases, the dimensionality of the prompt embedding space also increases Raffel et al. (2020); Meta (2023). This scaling leads to unnecessary complexity, as full dimensionality is often not required for task adaptation (Aghajanyan et al., 2021; Qin et al., 2021). Consequently, optimizing in this expanded space becomes inefficient in parameter’s usage and may also increase the risk of overfitting, especially for less complex tasks or with limited training data.
@@
In this paper, we propose Ultra-Low-Dimensional Prompt Tuning (ULPT), a novel approach that decouples the prompt and model dimensions, which enables to learn prompt embeddings in an ultra-low-dimensional space. A naïve attempt is to jointly optimize the ultra-low-dimensional embeddings with an up-projection matrix (Xiao et al., 2023; Guo et al., 2024), but the learnable up-projection matrix may result in more trainable parameters than vanilla prompt tuning. We avoid this overhead by employing a random but frozen matrix for the up-projection, as shown in Figure 1a. We further introduce lightweight, learnable shift and scale embedding vectors to better align the randomly projected embeddings with the model’s prompt space (Wu et al., 2024c).
We provide a convergence analysis for ULPT, and further show that a low-dimensional space with random projection can effectively approximate high-rank information that preserve the relational structure of embeddings, which is crucial for attention mechanisms in LLMs that depend on pairwise dot products between embeddings Vaswani et al. (2017).
@@
Figure 1: Overview of our approach. (a) ULPT up-projects ultra-low-dimensional embeddings with a random but fixed matrix. (b) ULPT can significantly reduce parameters storage for LLMs customization.
In addition, the random projection into ultra-low dimensions introduces a controllable tradeoff between prompt dimension and length under a fixed parameter budget. We empirically demonstrate that allocating more tokens with lower-dimensional embeddings yields greater expressivity than using fewer high-dimensional tokens. This makes ULPT well-suited for massive LLM customization, such as per-user tuning while keeping storage footprints minimal, as shown in Figure 1b.
We evaluated ULPT across over 20 NLP tasks, including GLUE Wang et al. (2018) and SuperGLUE Wang et al. (2019) for language understanding, MRQA Fisch et al. (2019) for question answering, GSM8K Cobbe et al. (2021) and MBPP Austin et al. (2021) for complex reasoning, as well as four additional tasks covering commonsense reasoning. The results demonstrate that prompt tuning via ultra-low-dimensional optimization matches or surpasses the performance of fully parameterized prompt tuning while saving up to 98% of trainable parameters. With an appropriate dimension, ULPT outperforms recent parameter-efficient fine-tuning methods, while requiring much fewer trainable parameters.
@@
  • We introduce ULPT, which optimizes prompts in a low-dimensional space with a random up-projection, drastically reducing trainable parameters while maintaining performance.
  • Theoretically, we show that ULPT effectively approximates high-rank structures, which preserves embedding relational structures that are essential for attention mechanisms in LLMs.
  • Empirically, we demonstrate that ULPT matches or surpasses vanilla prompt tuning across over 20 NLP tasks while saving trainable parameters by up to 98%. Scaling to higher dimensions for optimization, it outperforms recent efficient tuning methods with much fewer trainable parameters.
@@ 2 Related Work
With the rapid growth of pretrained neural networks, researchers have investigated parameter-efficient fine-tuning methods that update only a small set of parameters while maintaining high performance. One straightforward way is to tune specific components of the model. For example, BitFit updates only the bias terms Ben Zaken et al. (2022), and LayerNorm tuning only trains the layer-norm parameters Zhao et al. (2024). Another line of work involves introducing and training small, task-specific non-linear modules, such as Adapters Houlsby et al. (2019) and AdapterDrop Rücklé et al. (2021). Other methods steer the activation representations either globally Wu et al. (2024b); Pan et al. (2024) or locally Yin et al. (2024).
@@
Two prominent paradigms are low-rank adaptation (LoRA; Hu et al., 2022) and prompt tuning methods Lester et al. (2021), which are more related to our work. They will be further elaborated below.
Low-rank adaptation.
@@
Hu et al. (2022) assume that weight updates can be approximated by low-rank matrices and propose a low-rank adaptation (LoRA) method for fine-tuning a model. Building upon this foundational work, many extensions have been developed to enhance LoRA’s performance. For example, ReLoRA Lialin et al. (2024) iteratively trains and merges low-rank adapters to achieve high-rank updates. Hayou et al. (2024) propose learning low-rank matrices with different learning rates. Wu et al. (2024a) explore training a mixture of LoRA modules and leverage dynamic routing mechanisms for different task distributions or domains.
However, for large models, LoRA still requires a considerable number of trainable parameters, hindering efficient storage of task adaptations. To address this limitation, several studies have explored ways to further improve parameter efficiency. For example, VeRA Kopiczko et al. (2024) approximates LoRA weight update by using random matrices combined with two trainable scaling vectors. FourierFT Gao et al. (2024) learns a sparse set of spectral coefficients in the frequency domain and reconstructing weight updates via inverse Fourier transform. While these methods fine-tune in the weight space, our ULPT operates in the input embedding space and achieves better performance with lower training memory.
Prompt tuning.
@@
Shin et al. (2020) introduce the concept of learning prompt tokens to elicit knowledge from LLMs. Subsequently, Lester et al. (2021) extend this idea to continuous prompt tuning, where prompt embeddings are optimized through gradient descent. Building on this, Shi and Lipani (2024) observe that redistributing parameters to learn offsets for input token embeddings can enhance performance. Lan et al. (2025) decompose prompt embeddings using SVD for more meaningful initialization. In parallel, multi-task prompt tuning has been explored, where the learned prompt parameters are reused across different tasks Wang et al. (2023). Closely related to our work, Xiao et al. (2023) decompose the prompt embedding matrix into two low-rank components: a low-dimensional prompt matrix and a learnable up-projection matrix. By contrast, our ULPT pushes parameter reduction further by using frozen and random up-projection matrix, which drastically lowers the number of trainable parameters while preserving performance. Our method is supported by random-projection theory Bingham and Mannila (2001) and its recent applications in gradient compression Hao et al. (2024).
@@ 3.1 Problem Formulation
Figure 2: Distribution of prompt embedding values over 100 prompt tokens. We randomly selected $20$ dimensions from the original prompt embeddings, which have 768 dimensions as in the T5-base model.
3.2 Ultra-Low-Dimensional Prompt Tuning
@@
A naïve implementation of this decomposition, as in DPT Xiao et al. (2023), treats both $\bm{Z}$} and $\bm{P}$} as learnable parameters. This reduces the number of trainable parameters to $nr+rd$. However, the $rd$ term quickly becomes dominant as the model dimension $d$ grows or when separate $r\times d$ matrices must be maintained for multiple tasks, making DPT scale poorly in both parameter count and storage cost.
@@
In this way, we eliminate the need for storing the up-project matrix entirely, reducing the learnable parameters from $nr+rd$ to $nr$ (plus one extra random seed). Empirically, we find that freezing $\tilde{{\bm{P}}}$ mitigates overfitting, particularly when fine-tuning on small datasets.
In our pilot study, we observe that typical prompt embeddings $\bm{E}$}, even without low-rank treatment, exhibit significant variation across different dimensions, as shown in Figure 2. For comparison, we report the distribution of pretrained embeddings in Appendix B.1, showing that they exhibit much smaller variation than the learned task-specific prompt embeddings. These variations may hinder effective training, therefore, we further introduce a learnable shift embedding $\bm{s}\in\mathbb{R}^{d}$ and a learnable scale embedding $\bm{b}\in\mathbb{R}^{d}$ to adjust the projected embeddings to ensure better alignment with the varying distributions across dimensions. Notice that the shift and scale embeddings are shared across different prompt token positions, but may vary for different tasks.
@@
where $z_{ik}$ and $\tilde{p}_{kj}$ are an entry in $\bm{Z}$} and $\tilde{\bm{P}}$ matrices, respectively; $s_{j}$ and $b_{j}$ are an entry in $\bm{s}$ and $\bm{b}$ vectors, respectively.
@@ 3.3 Theoretical Analyses
theoremexpr Let ${\bm{e}}_{1},\dots,{\bm{e}}_{n}\in\mathbb{R}^{d}$ be the embedding vectors in a high-dimensional space. Let ${\bm{P}}\in\mathbb{R}^{r\times d}$ be a random projection matrix where each element $p_{i,j}\sim\mathcal{N}(0,1/r)$, and let ${\bm{z}}_{i}={\bm{P}}{\bm{e}}_{i}\in\mathbb{R}^{r}$ be the projected low-dimensional vectors.
Then, with confidence at least $1-\delta$, we have
@@
for all $i,j\in[n]$, given that $r\geq C\epsilon^{-2}\log(2n/\delta)$ for a sufficiently large constant $C$.
@@
In essence, our theorem asserts that by projecting data using a random matrix $\bm{P}$}, the pairwise $L^{2}$ distances of the original high-dimensional vectors are preserved for all $(i,j)$ pairs with high probability. Crucially, the projected dimension $r$ scales logarithmically with the number of embeddings $n$, rather than the original dimension $d$, demonstrating a favorable property of scaling.
It should be noted that, although our theorem uses $L^{2}$ as the metric, it can be extended to the dot-product metric by $\|{\bm{x}}-{\bm{y}}\|^{2}=\|{\bm{x}}\|^{2}+\|{\bm{y}}\|^{2}-2{\bm{x}}\cdot{\bm{y}}$. Practically, since LLMs compute attention using pairwise dot products between embeddings Vaswani et al. (2017), Theorem 1 implies that our up-projected low-dimensional prompts preserve the relational structure of full-dimensional embeddings.
@@
The above theorem shows the expressiveness of the low-dimensional space. We assert in the following theorem that, given a random up-projection matrix, the optimal low-dimensional embeddings can be learned by gradient descent under mild assumptions.
@@
Theorem 3.3 shows that, even with the naïve gradient descent, the fixed random matrix $\bm{P}$} does not hinder the optimization procedure. By combining Theorem 1, we theoretically justify our overall practice of ULPT.